# Mutation-Guided Unit Test Generation with a Large Language Model

### v4

@@ VI Discussion
Finally, although MutGen achieves high mutation scores on the selected subjects, there remain open challenges in pushing these scores even higher. For example, guiding LLMs to effectively kill mutants produced by certain mutation operators, such as VoidMethodCalls, remains difficult. In our evaluation on HumanEval-Java, this operator had only a 69.2% killing ratio. Developing strategies to better handle such challenging mutants is an important direction for future research.
@@ VII Conclusion
In this paper, we propose MutGen, a mutation-guided, LLM-based approach for automatically generating unit tests with high fault detection capability. The core innovation of MutGen lies in incorporating mutation feedback into the prompt to help LLMs generate tests that kill more mutants. In addition, two components—code summarization and fixing—further enhance the effectiveness of the approach.
@@
Moreover, MutGen employs an iterative generation process to push the limits of LLMs in killing mutants. We further analyze the causes of remaining live and uncovered mutants, as well as the impact of different mutation operators, offering insights and directions for future research.

### v5

@@ VI Discussion
Finally, although MutGen achieves high mutation scores on the selected subjects, there remain open challenges in further improving these scores. For example, guiding LLMs to effectively kill mutants produced by certain mutation operators, such as VoidMethodCalls, remains difficult. In our evaluation on HumanEval-Java, this operator had only a 69.2% killing ratio. Developing strategies to better handle such challenging mutants is an important direction for future research. We also observe that some generated test cases that are not repaired by the fixing stage are close to killing hard-to-detect mutants, but fail due to issues such as incorrect assertions or incomplete oracles. Since our focus here is on test generation, addressing the impact of such unfixed test cases is left as future work, where more advanced repair strategies may further improve mutation scores. In addition, an important avenue for future work is to more efficiently characterize the upper bound of LLM mutation scores across diverse target programs. In particular, identifying how quickly an LLM approaches its performance limit and how this limit varies across program sizes, mutation operators, and code structures remains an open problem.
@@ VII Conclusion
In this paper, we propose MutGen, a mutation-guided, LLM-based approach for automatically generating unit tests with high fault detection capability. The core innovation of MutGen lies in incorporating mutation feedback into the prompt to help LLMs generate tests that kill more mutants. In addition, two components—code summarization and code fixing—further enhance the approach’s effectiveness.
@@
Moreover, MutGen employs an iterative generation process to push the limits of LLMs in killing mutants. We further analyze the causes of remaining live and uncovered mutants and the impact of different mutation operators, offering insights and directions for future research.